OpenStudy (anonymous):

Solve the absolute value inequality: |2x - 3| < 5 x < -1 and x > 4 x > -1 and x < 4 x < -1 or x > 4 x > -4 and x < 4

OpenStudy (anonymous):

Where is the graph?

OpenStudy (shamil98):

you have two equations. 2x - 3 < 5 and -(2x-3) < 5 or simply.. 2x - 3 > - 5

OpenStudy (shamil98):

Just solve for x.

OpenStudy (anonymous):

If \[|x| < y\] then \[-y < x < y\] Therefore, \[|2x-3| < 5\]\[= -5 < 2x-3 < 5\] Now solve for x.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Isolate it. First you're going to add 3 to all sides of the inequality to remove the -3 from the center. \[-5+3 < 2x-3+3 < 5+3\]\[=-2<2x<8\]Now you have to divide both sides by 2. \[\frac{-2}{2} < \frac{2x}{2} < \frac{8}{2}\]\[=-1<x<4\] Therefore, \[x > -1\] and \[x < 4\]